# The Active and Noise-Tolerant Strategic Perceptron

### v1

@@ Abstract
Our main result is an algorithm for actively learning linear separators in the strategic setting that preserves the exponential improvement in label complexity over passive learning previously obtained only in the non-strategic case. Specifically, for data drawn uniformly from the unit sphere, we show that a modified version of the Active Perceptron algorithm (Dasgupta2005Analysis; DBLP:conf/nips/YanZ17) achieves excess error $\varepsilon$ using only $\tilde{O}\!\left(d\ln\tfrac{1}{\varepsilon}\right)$ label queries and incurs at most $\tilde{O}\!\left(d\ln\tfrac{1}{\varepsilon}\right)$ additional mistakes relative to the optimal classifier, even in the nonrealizable case, when a $\tilde{\Omega}(\varepsilon)$ fraction of inputs have inconsistent labels with the optimal classifier. The algorithm is computationally efficient and, under these distributional assumptions, requires substantially fewer label queries than prior work on strategic Perceptron (DBLP:conf/sigecom/AhmadiBBN21).
@@ 1 Introduction
Active learning is a well-established framework within machine learning that addresses the problem of efficiently training a model when labeled data is costly or limited Dasgupta2005Analysis; Balcan2014. Rather than assuming the learner has access to all labels, the learner selectively queries labels for the most informative examples. Consider a hiring, admissions, or loan-approval setting where evaluating each candidate requires careful human review. Because each label is expensive, the learner must decide which applications to examine in order to make the most progress in improving model accuracy. Active learning methods formalize this process: they use the features of each example to estimate whether obtaining its label will meaningfully reduce uncertainty and improve the model’s predictions. The main challenge of active learning is to design selection rules that allow the learner to identify these high-value examples while keeping the total labeling cost low.
@@
Strategic classification, on the other hand, is an emerging area of modern machine learning research that models scenarios where input features are provided by individuals who might manipulate them to receive better outcomes (DBLP:conf/kdd/BrucknerS11; DBLP:conf/innovations/HardtMPW16). This phenomenon arises in many decision-making systems where individuals are evaluated based on observable features. Returning to the same examples, suppose a classifier determines whether someone is shortlisted for an interview, admitted to a program, or offered a favorable loan rate. Individuals may modify their observable features—such as highlighting particular achievements, restructuring their resume, or inflating certain metrics—to increase the likelihood of being classified positively. These adjustments may not reflect genuine improvements in qualification or creditworthiness but are instead strategic changes made in response to the classifier. The main challenge in strategic classification is that the learner only observes manipulated features. This issue is especially critical in an online setting where data points arrive sequentially and the learner refines the classification rule over time: individuals react to the current classifier, causing the manipulation behavior to transform—even for similar underlying cases—and the learner’s updates rely on data already distorted as a result of manipulation in response to past classifiers. Finally, manipulation behavior can be discontinuous: if the cost of manipulation is just below the perceived benefit of being classified as positive, an individual will manipulate; if just above, they will not. As a result, standard online learning algorithms that achieve bounded mistakes in non-strategic settings may cycle indefinitely and make unbounded errors, even when a perfect classifier exists DBLP:conf/sigecom/AhmadiBBN21.
@@
In this work, we initiate the study of active learning for strategic classification. The main motivation for combining these two areas is that many real-world settings—including the aforementioned scenarios such as hiring, admissions, or loan-approval—must contend both with strategically altered data and costly expert labeling procedures. In contrast to prior work on strategic classification, which has largely focused on fully supervised settings [^1] An independent concurrent work is zhao2025online that studies strategic classification with partial feedback. where human experts label all training examples, our objective is to design learning algorithms that take a more active role than in the classic fully supervised setting, enabling learning with far fewer label requests. Beyond the individual challenges of active learning and strategic classification, we face unique difficulties arising from their combination. Because we operate in a strategic setting in which features may have been manipulated, there is a risk that the learner requests labels for examples that are ultimately uninformative, derailing the active learning process and leading to classifiers that perform poorly under the original data.
@@
We study an online linear classification problem in which the individuals being classified are strategic and the labels are expensive to obtain. Following the convention in strategic classification DBLP:conf/sigecom/AhmadiBBN21, each individual arriving at the classifier wishes to be classified positively and, if necessary, will manipulate their feature vector to achieve this outcome. More formally, an individual’s true feature vector is $\bm{z}$, but they may choose to report a manipulated vector $\bm{x}$ if it results in receiving a positive classification with the current classification rule. The manipulation comes at a cost, which reflects how far their reported vector $\bm{x}$ is from their true vector $\bm{z}$. Specifically, we model individuals as utility-maximizing agents, where the utility is defined as the value received from the classification outcome minus the cost of manipulation. The goal of each individual is to optimize $\bm{x}$ to maximize $\text{value}(\bm{x})-\text{cost}(\bm{z},\bm{x})]$, where $\text{value}(\bm{x})=1$ if the reported vector $\bm{x}$ is classified as positive and 0 if classified as negative. The cost function $\text{cost}(\bm{z},\bm{x})$ quantifies the cost of manipulating the features from $\bm{z}$ to $\bm{x}$. In this setting, if an individual can manipulate their features at a cost of at most 1 to change their classification from negative to positive, they will do so in the least costly way; otherwise, they will not manipulate their features. In line with standard active-learning assumptions (Dasgupta2005Analysis; DBLP:conf/colt/BalcanBZ07; DBLP:conf/icml/BalcanBL06; Balcan2014; hanneke2014active), we assume the feature vectors $\bm{z}$ are drawn from an underlying distribution. Formally, we assume that the true feature vectors $\bm{z}$ of individuals are uniformly distributed within a $d$-dimensional unit ball centered at the origin. In the realizable case, there exists a true classifier $\bm{u}$ such that $\bm{u}\cdot\bm{z}\geq 0$ for all positively labeled points and $\bm{u}\cdot\bm{z}<0$ for all negatively labeled points. In the nonrealizable case, a certain proportion of points may have labels inconsistent with the best homogeneous linear classifier $\bm{u}$, reflecting noise in the data. The task for the learning algorithm is to overcome the manipulations and accurately classify the true features, $\bm{z}$, while minimizing the number of labeling queries.
@@
We use the non-strategic active learning algorithms of DBLP:conf/nips/YanZ17 as our baseline, and generalize the prediction, label-query, and update rules in several aspects. A key adaptation in our algorithm is to focus label queries on unmanipulated examples: we request labels only for points classified as negative and lying in the active learning query region, since utility-maximizing agents do not manipulate when labeled negative, these samples are reliable. Furthermore, symmetry in misclassification ensures that ignoring misclassified positives is harmless: for hyperplanes through the origin, each positive has a corresponding negative with opposite coordinates, and both provide identical directional information for Perceptron-style updates. Consequently, concentrating solely on misclassified negatives preserves the theoretical guarantees while ensuring all queried data reflect true features. Another modification compared to prior work in active learning is raising the threshold for positive classification, as in DBLP:conf/sigecom/AhmadiBBN21: by requiring a strictly positive dot product with the weight vector, truly positive points eventually lie on the correct side (or can strategically manipulate to reach it), while negatives face prohibitively high manipulation costs. Although applied successfully in the realizable setting, the positive threshold could not be paired with prior update rules used in the strategic setting to accommodate the nonrealizable case. Our analysis includes a reduction from the strategic case to the non-strategic case.
@@
    Noise in Strategic Classification: In the nonrealizable case, we achieve excess error $\Theta(\varepsilon)$ after requesting only $\tilde{O}\left(d\ln\frac{1}{\varepsilon}\right)$ labels and making an additive $\tilde{O}\left(d\ln\frac{1}{\varepsilon}\right)$ mistakes compared to the best classifier, when the $\tilde{\Omega}(\varepsilon)$ fraction of the inputs are inverted. We resolve an open problem posed by DBLP:conf/sigecom/AhmadiBBN21 regarding handling noise in classification, moving beyond perfect separability. Previous techniques were insufficient for addressing this issue.
@@ 1.1 Related Literature
Active learning has a long-standing history in machine learning (Balcan2014). The central idea is that a learner can achieve better generalization with fewer labeled examples by selectively querying the most informative ones. DBLP:series/synthesis/2012Settles provides an extensive survey of active learning algorithms and their applications, highlighting the potential efficiency gains of this approach. From a theoretical perspective, key paradigms and analysis frameworks include disagreement-based active learning, first studied in the presence of noise by DBLP:conf/icml/BalcanBL06, and further developed by many others (DBLP:conf/nips/DasguptaHM07; DBLP:journals/jmlr/Koltchinskii10; DBLP:conf/nips/BeygelzimerHLZ10; DBLP:conf/icml/Hanneke07). Another widely studied and more practical paradigm is margin-based active learning, where the algorithm queries only examples near the current decision boundary. Our work falls into this category (DBLP:conf/nips/YanZ17; Dasgupta2005Analysis; DBLP:conf/colt/BalcanBZ07; DBLP:conf/colt/BalcanL13; DBLP:conf/stoc/AwasthiBL14), and is most closely related to DBLP:conf/nips/YanZ17, as discussed in Section 3.1.
@@
The study of strategic classification has gained increasing attention in recent years, motivated by the need to understand how individuals or entities may “game” machine learning systems. The main application domains are admissions, hiring, and financial decision-making (e.g., loan lending), where individuals have incentives to modify their input data. DBLP:conf/innovations/HardtMPW16 introduced foundational models of strategic classification, in which individuals manipulate their feature vectors to obtain more favorable outcomes. This line of work has since been extended to examine how classifiers can be designed to be robust to such manipulations DBLP:journals/jmlr/BrucknerKS12; DBLP:conf/fat/MilliMDH19. Several other works explore variations of the strategic classification model, including DBLP:conf/sigecom/DongRSWW18; DBLP:conf/forc/BravermanG20; DBLP:conf/nips/HarrisPW23, but all rely on fully labeled training data. In the application domains of strategic classification, acquiring labeled data—required in all prior work—is often costly, as it typically involves human expert judgment. This highlights the importance of developing algorithms that can learn effectively with fewer labeled examples.
@@
Prior work on online binary ($\pm 1$) classification in strategic settings—aimed at optimizing classification accuracy—has primarily focused on two cases: (i) when the original data is perfectly separable, as in DBLP:conf/sigecom/AhmadiBBN21, who showed that guarantees achievable in this setting can fail under even slight inseparability (noise); and (ii) when the data is not separable, but the algorithm’s performance degrades arbitrarily with the level of noise, as in DBLP:conf/nips/Chen0P20. In both cases, prior work falls short of providing robust guarantees in the presence of moderate noise, especially while maintaining low label complexity.
@@
The literatures on strategic classification and active learning both require certain modeling assumptions to obtain theoretical guarantees. Our model for strategic agents follows the framework of DBLP:conf/sigecom/AhmadiBBN21, but we study it in a more challenging active learning setting where the learner selectively queries labels to minimize the need for human intervention. Even in the simpler non-strategic case, active learning requires distributional assumptions to provably achieve label-complexity improvements over passive learning (Dasgupta2005Analysis; DBLP:conf/colt/BalcanBZ07; DBLP:conf/icml/BalcanBL06; Balcan2014; hanneke2014active). For clarity of exposition, we adopt one of the most widely used such assumptions for learning linear separators: the true feature vectors are drawn uniformly at random. In Section 4, we further explain why our results do not rely critically on this assumption.
@@
The feature data distortion caused by strategic manipulation is fundamentally different from the feature distortion considered in other classic learning theory models such as malicious-noise model valiant1984theory; DBLP:conf/stoc/KearnsL88; JMLR:v10:klivans09a; DBLP:conf/stoc/AwasthiBL14; DBLP:journals/jacm/AwasthiBL17. Although in both settings the observed data is a distorted version of the underlying distribution, the nature of this distortion differs significantly. The malicious-noise model relies critically on the assumption that only a small fraction of examples are corrupted arbitrarily, while the remaining examples are drawn cleanly from the true distribution. In contrast, in strategic settings every agent is strategic, and their observed features may be manipulated whenever doing so benefits them. Since arbitrary manipulation yields no meaningful theoretical guarantees, the literature on strategic classification typically assumes that manipulation is structured: agents adjust their features to optimize their utility in response to the current classifier. Under this assumption, the resulting distortion is distribution-dependent and tightly coupled with the learner’s current classification rule. Algorithms designed for malicious noise, however, rely heavily on having a large fraction of uncorrupted examples in each round—a property that strategic settings do not guarantee. Consequently, because the structure of distortion differ so substantially between malicious noise and strategic classification, it is unclear whether results from one model should transfer to the other. One of our contributions is to show that a style of localization algorithms DBLP:conf/stoc/AwasthiBL14; DBLP:journals/jacm/AwasthiBL17 developed for malicious-noise tolerance can, in fact, be adapted to strategic classification.
@@
The intersection of active learning and strategic classification is a relatively new area of research. While most active learning models assume access to truthful data, strategic classification explicitly accounts for agents who may manipulate their features. From a theoretical perspective, a key motivation for studying this combination is that ideas from active learning have historically been transformative in improving passive learning algorithms. In several challenging learning problems where existing passive methods failed to obtain optimal guarantees, incorporating active learning techniques has led to major breakthroughs (DBLP:conf/stoc/AwasthiBL14; DBLP:conf/colt/AwasthiBHU15; DBLP:conf/colt/AwasthiBHZ16; DBLP:journals/jacm/AwasthiBL17). In this work, we show that active learning principles can likewise overcome fundamental limitations in classifying strategic entities. In particular, while prior passive algorithms struggled in this setting, selectively querying labels and focusing on informative examples yields stronger theoretical guarantees and more robust performance. The central challenge we address is how to incorporate active learning techniques when interacting with strategic agents who may manipulate their data, thereby balancing the need for efficient learning with robustness to strategic behavior.

### v2

@@ Abstract
Our main result is an algorithm for actively learning linear separators in the strategic setting that preserves the exponential improvement in label complexity over passive learning previously obtained only in the non-strategic case. Specifically, for data drawn uniformly from the unit sphere, we show that a modified version of the Active Perceptron algorithm (DKM, 05; YZ, 17) achieves excess error $\varepsilon$ using only $\tilde{O}\!\left(d\ln\tfrac{1}{\varepsilon}\right)$ label queries and incurs at most $\tilde{O}\!\left(d\ln\tfrac{1}{\varepsilon}\right)$ additional mistakes relative to the optimal classifier, even in the nonrealizable case, when a $\tilde{\Omega}(\varepsilon)$ fraction of inputs have inconsistent labels with the optimal classifier. The algorithm is computationally efficient and, under these distributional assumptions, requires substantially fewer label queries than prior work on strategic Perceptron (ABBN, 21).
@@ 1 Introduction
Active learning is a well-established framework within machine learning that addresses the problem of efficiently training a model when labeled data is costly or limited DKM (05); BU (14). Rather than assuming the learner has access to all labels, the learner selectively queries labels for the most informative examples. Consider a hiring, admissions, or loan-approval setting where evaluating each candidate requires careful human review. Because each label is expensive, the learner must decide which applications to examine in order to make the most progress in improving model accuracy. Active learning methods formalize this process: they use the features of each example to estimate whether obtaining its label will meaningfully reduce uncertainty and improve the model’s predictions. The main challenge of active learning is to design selection rules that allow the learner to identify these high-value examples while keeping the total labeling cost low.
@@
Strategic classification, on the other hand, is an emerging area of modern machine learning research that models scenarios where input features are provided by individuals who might manipulate them to receive better outcomes (BS, 11; HMPW, 16). This phenomenon arises in many decision-making systems where individuals are evaluated based on observable features. Returning to the same examples, suppose a classifier determines whether someone is shortlisted for an interview, admitted to a program, or offered a favorable loan rate. Individuals may modify their observable features—such as highlighting particular achievements, restructuring their resume, or inflating certain metrics—to increase the likelihood of being classified positively. These adjustments may not reflect genuine improvements in qualification or creditworthiness but are instead strategic changes made in response to the classifier. The main challenge in strategic classification is that the learner only observes manipulated features. This issue is especially critical in an online setting where data points arrive sequentially and the learner refines the classification rule over time: individuals react to the current classifier, causing the manipulation behavior to transform—even for similar underlying cases—and the learner’s updates rely on data already distorted as a result of manipulation in response to past classifiers. Finally, manipulation behavior can be discontinuous: if the cost of manipulation is just below the perceived benefit of being classified as positive, an individual will manipulate; if just above, they will not. As a result, standard online learning algorithms that achieve bounded mistakes in non-strategic settings may cycle indefinitely and make unbounded errors, even when a perfect classifier exists ABBN (21).
@@
In this work, we initiate the study of active learning for strategic classification. The main motivation for combining these two areas is that many real-world settings—including the aforementioned scenarios such as hiring, admissions, or loan-approval—must contend both with strategically altered data and costly expert labeling procedures. In contrast to prior work on strategic classification, which has largely focused on fully supervised settings [^1] An independent concurrent work is ZML (25) that studies strategic classification with partial feedback. where human experts label all training examples, our objective is to design learning algorithms that take a more active role than in the classic fully supervised setting, enabling learning with far fewer label requests. Beyond the individual challenges of active learning and strategic classification, we face unique difficulties arising from their combination. Because we operate in a strategic setting in which features may have been manipulated, there is a risk that the learner requests labels for examples that are ultimately uninformative, derailing the active learning process and leading to classifiers that perform poorly under the original data.
@@
We study an online linear classification problem in which the individuals being classified are strategic and the labels are expensive to obtain. Following the convention in strategic classification ABBN (21), each individual arriving at the classifier wishes to be classified positively and, if necessary, will manipulate their feature vector to achieve this outcome. More formally, an individual’s true feature vector is $\bm{z}$, but they may choose to report a manipulated vector $\bm{x}$ if it results in receiving a positive classification with the current classification rule. The manipulation comes at a cost, which reflects how far their reported vector $\bm{x}$ is from their true vector $\bm{z}$. Specifically, we model individuals as utility-maximizing agents, where the utility is defined as the value received from the classification outcome minus the cost of manipulation. The goal of each individual is to optimize $\bm{x}$ to maximize $\text{value}(\bm{x})-\text{cost}(\bm{z},\bm{x})]$, where $\text{value}(\bm{x})=1$ if the reported vector $\bm{x}$ is classified as positive and 0 if classified as negative. The cost function $\text{cost}(\bm{z},\bm{x})$ quantifies the cost of manipulating the features from $\bm{z}$ to $\bm{x}$. In this setting, if an individual can manipulate their features at a cost of at most 1 to change their classification from negative to positive, they will do so in the least costly way; otherwise, they will not manipulate their features. In line with standard active-learning assumptions (DKM, 05; BBZ, 07; BBL, 06; BU, 14; Han, 14), we assume the feature vectors $\bm{z}$ are drawn from an underlying distribution. Formally, we assume that the true feature vectors $\bm{z}$ of individuals are uniformly distributed within a $d$-dimensional unit ball centered at the origin. In the realizable case, there exists a true classifier $\bm{u}$ such that $\bm{u}\cdot\bm{z}\geq 0$ for all positively labeled points and $\bm{u}\cdot\bm{z}<0$ for all negatively labeled points. In the nonrealizable case, a certain proportion of points may have labels inconsistent with the best homogeneous linear classifier $\bm{u}$, reflecting noise in the data. The task for the learning algorithm is to overcome the manipulations and accurately classify the true features, $\bm{z}$, while minimizing the number of labeling queries.
@@
We use the non-strategic active learning algorithms of YZ (17) as our baseline, and generalize the prediction, label-query, and update rules in several aspects. A key adaptation in our algorithm is to focus label queries on unmanipulated examples: we request labels only for points classified as negative and lying in the active learning query region, since utility-maximizing agents do not manipulate when labeled negative, these samples are reliable. Furthermore, symmetry in misclassification ensures that ignoring misclassified positives is harmless: for hyperplanes through the origin, each positive has a corresponding negative with opposite coordinates, and both provide identical directional information for Perceptron-style updates. Consequently, concentrating solely on misclassified negatives preserves the theoretical guarantees while ensuring all queried data reflect true features. Another modification compared to prior work in active learning is raising the threshold for positive classification, as in ABBN (21): by requiring a strictly positive dot product with the weight vector, truly positive points eventually lie on the correct side (or can strategically manipulate to reach it), while negatives face prohibitively high manipulation costs. Although applied successfully in the realizable setting, the positive threshold could not be paired with prior update rules used in the strategic setting to accommodate the nonrealizable case. Our analysis includes a reduction from the strategic case to the non-strategic case.
@@
    Noise in Strategic Classification: In the nonrealizable case, we achieve excess error $\Theta(\varepsilon)$ after requesting only $\tilde{O}\left(d\ln\frac{1}{\varepsilon}\right)$ labels and making an additive $\tilde{O}\left(d\ln\frac{1}{\varepsilon}\right)$ mistakes compared to the best classifier, when the $\tilde{\Omega}(\varepsilon)$ fraction of the inputs are inverted. We resolve an open problem posed by ABBN (21) regarding handling noise in classification, moving beyond perfect separability. Previous techniques were insufficient for addressing this issue.
@@ 1.1 Related Literature
Active learning has a long-standing history in machine learning (BU, 14). The central idea is that a learner can achieve better generalization with fewer labeled examples by selectively querying the most informative ones. Set (12) provides an extensive survey of active learning algorithms and their applications, highlighting the potential efficiency gains of this approach. From a theoretical perspective, key paradigms and analysis frameworks include disagreement-based active learning, first studied in the presence of noise by BBL (06), and further developed by many others (DHM, 07; Kol, 10; BHLZ, 10; Han, 07). Another widely studied and more practical paradigm is margin-based active learning, where the algorithm queries only examples near the current decision boundary. Our work falls into this category (YZ, 17; DKM, 05; BBZ, 07; BL, 13; ABL, 14), and is most closely related to YZ (17), as discussed in Section 3.1.
@@
The study of strategic classification has gained increasing attention in recent years, motivated by the need to understand how individuals or entities may “game” machine learning systems. The main application domains are admissions, hiring, and financial decision-making (e.g., loan lending), where individuals have incentives to modify their input data. HMPW (16) introduced foundational models of strategic classification, in which individuals manipulate their feature vectors to obtain more favorable outcomes. This line of work has since been extended to examine how classifiers can be designed to be robust to such manipulations BKS (12); MMDH (19). Several other works explore variations of the strategic classification model, including DRS+ (18); BG (20); HPW (23), but all rely on fully labeled training data. In the application domains of strategic classification, acquiring labeled data—required in all prior work—is often costly, as it typically involves human expert judgment. This highlights the importance of developing algorithms that can learn effectively with fewer labeled examples.
@@
Prior work on online binary ($\pm 1$) classification in strategic settings—aimed at optimizing classification accuracy—has primarily focused on two cases: (i) when the original data is perfectly separable, as in ABBN (21), who showed that guarantees achievable in this setting can fail under even slight inseparability (noise); and (ii) when the data is not separable, but the algorithm’s performance degrades arbitrarily with the level of noise, as in CLP (20). In both cases, prior work falls short of providing robust guarantees in the presence of moderate noise, especially while maintaining low label complexity.
@@
The literatures on strategic classification and active learning both require certain modeling assumptions to obtain theoretical guarantees. Our model for strategic agents follows the framework of ABBN (21), but we study it in a more challenging active learning setting where the learner selectively queries labels to minimize the need for human intervention. Even in the simpler non-strategic case, active learning requires distributional assumptions to provably achieve label-complexity improvements over passive learning (DKM, 05; BBZ, 07; BBL, 06; BU, 14; Han, 14). For clarity of exposition, we adopt one of the most widely used such assumptions for learning linear separators: the true feature vectors are drawn uniformly at random. In Section 4, we further explain why our results do not rely critically on this assumption.
@@
The feature data distortion caused by strategic manipulation is fundamentally different from the feature distortion considered in other classic learning theory models such as malicious-noise model Val (84); KL (88); KLS (09); ABL (14, 17). Although in both settings the observed data is a distorted version of the underlying distribution, the nature of this distortion differs significantly. The malicious-noise model relies critically on the assumption that only a small fraction of examples are corrupted arbitrarily, while the remaining examples are drawn cleanly from the true distribution. In contrast, in strategic settings every agent is strategic, and their observed features may be manipulated whenever doing so benefits them. Since arbitrary manipulation yields no meaningful theoretical guarantees, the literature on strategic classification typically assumes that manipulation is structured: agents adjust their features to optimize their utility in response to the current classifier. Under this assumption, the resulting distortion is distribution-dependent and tightly coupled with the learner’s current classification rule. Algorithms designed for malicious noise, however, rely heavily on having a large fraction of uncorrupted examples in each round—a property that strategic settings do not guarantee. Consequently, because the structure of distortion differ so substantially between malicious noise and strategic classification, it is unclear whether results from one model should transfer to the other. One of our contributions is to show that a style of localization algorithms ABL (14, 17) developed for malicious-noise tolerance can, in fact, be adapted to strategic classification.
@@
The intersection of active learning and strategic classification is a relatively new area of research. While most active learning models assume access to truthful data, strategic classification explicitly accounts for agents who may manipulate their features. From a theoretical perspective, a key motivation for studying this combination is that ideas from active learning have historically been transformative in improving passive learning algorithms. In several challenging learning problems where existing passive methods failed to obtain optimal guarantees, incorporating active learning techniques has led to major breakthroughs (ABL, 14; ABHU, 15; ABHZ, 16; ABL, 17). In this work, we show that active learning principles can likewise overcome fundamental limitations in classifying strategic entities. In particular, while prior passive algorithms struggled in this setting, selectively querying labels and focusing on informative examples yields stronger theoretical guarantees and more robust performance. The central challenge we address is how to incorporate active learning techniques when interacting with strategic agents who may manipulate their data, thereby balancing the need for efficient learning with robustness to strategic behavior.